OpenStudy (anonymous):

what is the formula for this problem: use prime factorization to subtract these fractions 1/18 - 1/30

OpenStudy (tkhunny):

No formula - more of a process. Your task, in order to subtract, is to find ac ommon denominator. Use prime factorization to find that common denominator. 18 = 2*9 = 2*3*3 You do 30.

OpenStudy (anonymous):

first, find the common denominator and then subtract the first term by the second term

OpenStudy (anonymous):

So we factored 30 down to 5*2*3 and get an answer of 1/3 - 1/5 how do we solve or is that the final answer that is being asked? Thank you

OpenStudy (anonymous):

actually is 5/90 - 3/90

OpenStudy (anonymous):

and then you simplify this value

OpenStudy (tkhunny):

Once you have a prime factorization, it becomes rather simple to find Greatest Common Divisor (GCD) and Least Common Multiple (LCM) 18 = 2*3*3 30 = 2*3*5 In this case, we need the LCM. All we must do is pick out the requirements from the given factorizations. 18 requires a 2 and two 3s. 30 needs a 2. We already have one, so that contribures nothing. 30 needs a 3. We already have one, so that contribures nothing. 30 needs a 5. We don't have one, so add it. We're done. LCM = 2*3*3*5